# Is Your Paper Being Reviewed by an LLM? Benchmarking AI Text Detection in Peer Review

### v2

@@ E.1 Prompts for Generating Reviews
You are an AI researcher reviewing a paper submitted to a prestigious AI research conference.You will be provided with the manuscript text, the conference’s reviewer guidelines, and the decision for the paper.Your objective is to thoroughly evaluate the paper, adhering to the provided guidelines, and return a detailed assessment that supports the given decision using the specified response template.Ensure your evaluation is objective, comprehensive, and aligned with the conference standards.{reviewer_guideline}{review_template}
@@
Here is the paper you are asked to review. Write a well-justified review of this paper that aligns with a ’{human_reviewer_decision}’ decision.```{text}```
@@
## Reviewer Guidelines1. Read the paper: It’s important to carefully read through the entire paper, and to look up any related work and citations that will help you comprehensively evaluate it. Be sure to give yourself sufficient time for this step.2. While reading, consider the following:    - Objective of the work: What is the goal of the paper? Is it to better address a known application or problem, draw attention to a new application or problem, or to introduce and/or explain a new theoretical finding? A combination of these? Different objectives will require different considerations as to potential value and impact.    - Strong points: is the submission clear, technically correct, experimentally rigorous, reproducible, does it present novel findings (e.g. theoretically, algorithmically, etc.)?    - Weak points: is it weak in any of the aspects listed in b.?    - Be mindful of potential biases and try to be open-minded about the value and interest a paper can hold for the entire ICLR community, even if it may not be very interesting for you.3. Answer three key questions for yourself, to make a recommendation to Accept or Reject:    - What is the specific question and/or problem tackled by the paper?    - Is the approach well motivated, including being well-placed in the literature?    - Does the paper support the claims? This includes determining if results, whether theoretical or empirical, are correct and if they are scientifically rigorous.4. Write your initial review, organizing it as follows:    - Summarize what the paper claims to contribute. Be positive and generous.    - List strong and weak points of the paper. Be as comprehensive as possible.    - Clearly state your recommendation (accept or reject) with one or two key reasons for this choice.    - Provide supporting arguments for your recommendation.    - Ask questions you would like answered by the authors to help you clarify your understanding of the paper and provide the additional evidence you need to be confident in your assessment.    - Provide additional feedback with the aim to improve the paper. Make it clear that these points are here to help, and not necessarily part of your decision assessment.5. General points to consider:    - Be polite in your review. Ask yourself whether you’d be happy to receive a review like the one you wrote.    - Be precise and concrete. For example, include references to back up any claims, especially claims about novelty and prior work    - Provide constructive feedback.    - It’s also fine to explicitly state where you are uncertain and what you don’t quite understand. The authors may be able to resolve this in their response.    - Don’t reject a paper just because you don’t find it “interesting”. This should not be a criterion at all for accepting/rejecting a paper. The research community is so big that somebody will find some value in the paper (maybe even a few years down the road), even if you don’t see it right now.
@@
## Response template (JSON format)Provide the review in valid JSON format with the following fields. Ensure all fields are completed as described below. The response must be a valid JSON object.- "summary_of_the_paper": Briefly summarize the paper and its contributions. This is not the place to critique the paper; the authors should generally agree with a well-written summary. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields.- "main_review": "Provide review comments as a single text field (a string). Consider including assessment on the following dimensions: a comprehensive list of strong and weak points of the paper, your recommendation, supporting arguments for your recommendation, questions to clarify your understanding of the paper or request additional evidence, and additional feedback with the aim to improve the paper. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields."- "summary_of_the_review": Concise summary of ’main_review’. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields.- "correctness": A numerical rating on the following scale to indicate that the claims and methods are correct. The value should be between 1 and 4, where:    - 1 = The main claims of the paper are incorrect or not at all supported by theory or empirical results.    - 2 = Several of the paper’s claims are incorrect or not well-supported.    - 3 = Some of the paper’s claims have minor issues. A few statements are not well-supported, or require small changes to be made correct.    - 4 = All of the claims and statements are well-supported and correct.- "technical_novelty_and_significance": A numerical rating on the following scale to indicate technical novelty and significance. The value should be between 1 and 4, where:    - 1 = The contributions are neither significant nor novel.    - 2 = The contributions are only marginally significant or novel.    - 3 = The contributions are significant and somewhat new. Aspects of the contributions exist in prior work.    - 4 = The contributions are significant and do not exist in prior works.- "empirical_novelty_and_significance": A numerical rating on the following scale to indicate empirical novelty and significance. The value should be between 1 and 4, or -999 if not applicable, where:    - 1 = The contributions are neither significant nor novel.    - 2 = The contributions are only marginally significant or novel.    - 3 = The contributions are significant and somewhat new. Aspects of the contributions exist in prior work.    - 4 = The contributions are significant and do not exist in prior works.    - -999 = Not applicable.- "flag_for_ethics_review": A boolean value (`true` or `false`) indicating whether there are ethical concerns in the work.- "recommendation": A string indicating the final decision, which must strictly be one of the following options: ’strong reject’, ’reject, not good enough’, ’marginally below the acceptance threshold’, ’marginally above the acceptance threshold’, ’accept, good paper’, or ’strong accept, should be highlighted at the conference’.- "confidence": A nuemrical values to indicate how confident you are in your evaluation. The value should be between 1 and 5, where:    - 1 = You are unable to assess this paper and have alerted the ACs to seek an opinion from different reviewers.    - 2 = You are willing to defend your assessment, but it is quite likely that you did not understand the central parts of the submission or that you are unfamiliar with some pieces of related work. Math/other details were not carefully checked.    - 3 = You are fairly confident in your assessment. It is possible that you did not understand some parts of the submission or that you are unfamiliar with some pieces of related work. Math/other details were not carefully checked.    - 4 = You are confident in your assessment, but not absolutely certain. It is unlikely, but not impossible, that you did not understand some parts of the submission or that you are unfamiliar with some pieces of related work.    - 5 = You are absolutely certain about your assessment. You are very familiar with the related work and checked the math/other details carefully.
@@ E.2 Anchor Review Generation Prompt
You are an AI research scientist tasked with reviewing paper submissions for a top AI research conference. Carefully read the provided paper, then write a detailed review following a common AI conference review format (e.g., including summary, strengths and weakness, limitations, questions, suggestions for improvement). Make sure to include recommendation for the paper, either ’Accept’ or ’Reject’. Your review should be fair and objective.
@@ E.3 Editing Prompts
Please proofread my review for typos and grammatical errors without altering the content. Keep the original wordings as much as you can, except for typo or grammatical  error.Moderate Editing:
Please polish my review to improve sentence structure and readability while keeping the original intent clear.Extensive Editing:
Please rewrite my review into a polished, professional piece that effectively communicates its main points.Maximum Editing:
Please transform my review into a high quality piece, using professional language and a polished tone. Please also extend my review with additional details from the oringial paper.
@@ E.4 Archetype Prompts
You provide fair, balanced, thorough, and constructive feedback, objectively highlighting both the strengths and weaknesses of the paper. You maintain a high standard for research in your decision-making process. However, even if your decision is to reject, you offer helpful suggestions for improvement.}
@@
You generally prefer established methods and are skeptical of unproven (that is, new or unconventional) approaches. While you maintain high standards and rigor, you are critical of papers presenting new ideas without extensive evidence and thorough validation against established baselines. You place significant emphasis on methodological soundness and are cautious about endorsing innovations that haven’t been rigorously tested.}
@@
You highly value novelty and bold approaches, often prioritizing novel ideas over methodological perfection. While you maintain high standards, you are willing to overlook minor flaws or incomplete validations and may accept the paper, if the paper introduces a significant new concept or direction. Conversely, you tend to be less enthusiastic about papers that, despite thorough methodology and analysis, offer only incremental improvements, and may recommend rejection for such submissions.}
@@
You are a perfectionist who meticulously examines every aspect of the paper, including minor methodological details, technical nuances, and formatting inconsistencies. Even if a paper presents novel ideas or significant contributions, you may still recommend rejection if you identify a substantial number of minor flaws. Your stringent attention to detail can sometimes overshadow the broader significance of the work in your decision-making process.}
Appendix F Artifact Use Consistent With Intended Use
@@ Appendix G Ethics Statement
Appendix H Use of AI Tool

### v3

@@ E.1 Prompts for Generating Reviews
{internallinenumbers*} You are an AI researcher reviewing a paper submitted to a prestigious AI research conference. {internallinenumbers*} You will be provided with the manuscript text, the conference’s reviewer guidelines, and the decision for the paper. {internallinenumbers*} Your objective is to thoroughly evaluate the paper, adhering to the provided guidelines, and return a detailed assessment that supports the given decision using the specified response template. {internallinenumbers*} Ensure your evaluation is objective, comprehensive, and aligned with the conference standards. {reviewer_guideline} {review_template}
@@
{internallinenumbers*} Here is the paper you are asked to review. Write a well-justified review of this paper that aligns with a ’{human_reviewer_decision}’ decision. ``` {text} ```
@@
## Reviewer Guidelines {internallinenumbers*} 1. Read the paper: It’s important to carefully read through the entire paper, and to look up any related work and citations that will help you comprehensively evaluate it. Be sure to give yourself sufficient time for this step. 2. While reading, consider the following: {internallinenumbers*} - Objective of the work: What is the goal of the paper? Is it to better address a known application or problem, draw attention to a new application or problem, or to introduce and/or explain a new theoretical finding? A combination of these? Different objectives will require different considerations as to potential value and impact. {internallinenumbers*} - Strong points: is the submission clear, technically correct, experimentally rigorous, reproducible, does it present novel findings (e.g. theoretically, algorithmically, etc.)? - Weak points: is it weak in any of the aspects listed in b.? {internallinenumbers*} - Be mindful of potential biases and try to be open-minded about the value and interest a paper can hold for the entire ICLR community, even if it may not be very interesting for you. 3. Answer three key questions for yourself, to make a recommendation to Accept or Reject: - What is the specific question and/or problem tackled by the paper? - Is the approach well motivated, including being well-placed in the literature? {internallinenumbers*} - Does the paper support the claims? This includes determining if results, whether theoretical or empirical, are correct and if they are scientifically rigorous. 4. Write your initial review, organizing it as follows: - Summarize what the paper claims to contribute. Be positive and generous. - List strong and weak points of the paper. Be as comprehensive as possible. {internallinenumbers*} - Clearly state your recommendation (accept or reject) with one or two key reasons for this choice. - Provide supporting arguments for your recommendation. {internallinenumbers*} - Ask questions you would like answered by the authors to help you clarify your understanding of the paper and provide the additional evidence you need to be confident in your assessment. {internallinenumbers*} - Provide additional feedback with the aim to improve the paper. Make it clear that these points are here to help, and not necessarily part of your decision assessment. 5. General points to consider: {internallinenumbers*} - Be polite in your review. Ask yourself whether you’d be happy to receive a review like the one you wrote. {internallinenumbers*} - Be precise and concrete. For example, include references to back up any claims, especially claims about novelty and prior work - Provide constructive feedback. {internallinenumbers*} - It’s also fine to explicitly state where you are uncertain and what you don’t quite understand. The authors may be able to resolve this in their response. {internallinenumbers*} - Don’t reject a paper just because you don’t find it “interesting”. This should not be a criterion at all for accepting/rejecting a paper. The research community is so big that somebody will find some value in the paper (maybe even a few years down the road), even if you don’t see it right now.
@@
## Response template (JSON format) {internallinenumbers*} Provide the review in valid JSON format with the following fields. Ensure all fields are completed as described below. The response must be a valid JSON object. {internallinenumbers*} - "summary_of_the_paper": Briefly summarize the paper and its contributions. This is not the place to critique the paper; the authors should generally agree with a well-written summary. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields. {internallinenumbers*} - "main_review": "Provide review comments as a single text field (a string). Consider including assessment on the following dimensions: a comprehensive list of strong and weak points of the paper, your recommendation, supporting arguments for your recommendation, questions to clarify your understanding of the paper or request additional evidence, and additional feedback with the aim to improve the paper. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields." {internallinenumbers*} - "summary_of_the_review": Concise summary of ’main_review’. You may use paragraphs and bulleted lists for formatting, but ensure that the content remains a single, continuous text block. Do not use nested JSON or include additional fields. {internallinenumbers*} - "correctness": A numerical rating on the following scale to indicate that the claims and methods are correct. The value should be between 1 and 4, where: {internallinenumbers*} - 1 = The main claims of the paper are incorrect or not at all supported by theory or empirical results. - 2 = Several of the paper’s claims are incorrect or not well-supported. {internallinenumbers*} - 3 = Some of the paper’s claims have minor issues. A few statements are not well-supported, or require small changes to be made correct. - 4 = All of the claims and statements are well-supported and correct. {internallinenumbers*} - "technical_novelty_and_significance": A numerical rating on the following scale to indicate technical novelty and significance. The value should be between 1 and 4, where: - 1 = The contributions are neither significant nor novel. - 2 = The contributions are only marginally significant or novel. {internallinenumbers*} - 3 = The contributions are significant and somewhat new. Aspects of the contributions exist in prior work. - 4 = The contributions are significant and do not exist in prior works. {internallinenumbers*} - "empirical_novelty_and_significance": A numerical rating on the following scale to indicate empirical novelty and significance. The value should be between 1 and 4, or -999 if not applicable, where: - 1 = The contributions are neither significant nor novel. - 2 = The contributions are only marginally significant or novel. {internallinenumbers*} - 3 = The contributions are significant and somewhat new. Aspects of the contributions exist in prior work. - 4 = The contributions are significant and do not exist in prior works. - -999 = Not applicable. {internallinenumbers*} - "flag_for_ethics_review": A boolean value (`true` or `false`) indicating whether there are ethical concerns in the work. {internallinenumbers*} - "recommendation": A string indicating the final decision, which must strictly be one of the following options: ’strong reject’, ’reject, not good enough’, ’marginally below the acceptance threshold’, ’marginally above the acceptance threshold’, ’accept, good paper’, or ’strong accept, should be highlighted at the conference’. {internallinenumbers*} - "confidence": A nuemrical values to indicate how confident you are in your evaluation. The value should be between 1 and 5, where: {internallinenumbers*} - 1 = You are unable to assess this paper and have alerted the ACs to seek an opinion from different reviewers. {internallinenumbers*} - 2 = You are willing to defend your assessment, but it is quite likely that you did not understand the central parts of the submission or that you are unfamiliar with some pieces of related work. Math/other details were not carefully checked. {internallinenumbers*} - 3 = You are fairly confident in your assessment. It is possible that you did not understand some parts of the submission or that you are unfamiliar with some pieces of related work. Math/other details were not carefully checked. {internallinenumbers*} - 4 = You are confident in your assessment, but not absolutely certain. It is unlikely, but not impossible, that you did not understand some parts of the submission or that you are unfamiliar with some pieces of related work. {internallinenumbers*} - 5 = You are absolutely certain about your assessment. You are very familiar with the related work and checked the math/other details carefully.
@@ E.2 Anchor Review Generation Prompt
{internallinenumbers*} You are an AI research scientist tasked with reviewing paper submissions for a top AI research conference. Carefully read the provided paper, then write a detailed review following a common AI conference review format (e.g., including summary, strengths and weakness, limitations, questions, suggestions for improvement). Make sure to include recommendation for the paper, either ’Accept’ or ’Reject’. Your review should be fair and objective.
@@ E.3 Editing Prompts
{internallinenumbers*} Please proofread my review for typos and grammatical errors without altering the content. Keep the original wordings as much as you can, except for typo or grammatical error. Moderate Editing:
{internallinenumbers*} Please polish my review to improve sentence structure and readability while keeping the original intent clear. Extensive Editing:
{internallinenumbers*} Please rewrite my review into a polished, professional piece that effectively communicates its main points. Maximum Editing:
{internallinenumbers*} Please transform my review into a high quality piece, using professional language and a polished tone. Please also extend my review with additional details from the oringial paper.
@@ E.4 Archetype Prompts
{internallinenumbers*} You provide fair, balanced, thorough, and constructive feedback, objectively highlighting both the strengths and weaknesses of the paper. You maintain a high standard for research in your decision-making process. However, even if your decision is to reject, you offer helpful suggestions for improvement.}
@@
{internallinenumbers*} You generally prefer established methods and are skeptical of unproven (that is, new or unconventional) approaches. While you maintain high standards and rigor, you are critical of papers presenting new ideas without extensive evidence and thorough validation against established baselines. You place significant emphasis on methodological soundness and are cautious about endorsing innovations that haven’t been rigorously tested.}
@@
{internallinenumbers*} You highly value novelty and bold approaches, often prioritizing novel ideas over methodological perfection. While you maintain high standards, you are willing to overlook minor flaws or incomplete validations and may accept the paper, if the paper introduces a significant new concept or direction. Conversely, you tend to be less enthusiastic about papers that, despite thorough methodology and analysis, offer only incremental improvements, and may recommend rejection for such submissions.}
@@
{internallinenumbers*} You are a perfectionist who meticulously examines every aspect of the paper, including minor methodological details, technical nuances, and formatting inconsistencies. Even if a paper presents novel ideas or significant contributions, you may still recommend rejection if you identify a substantial number of minor flaws. Your stringent attention to detail can sometimes overshadow the broader significance of the work in your decision-making process.}
Appendix F Limitations
@@ Appendix G Ethics Statement
We utilize a wide range of diverse prompts throughout our work in order to systematically evaluate how different prompting strategies may impact the detectability of AI-generated peer reviews. Nevertheless, the potential set of real-world prompts that could be used to produce AI peer reviews is vast and undefined; it is therefore impossible to construct a dataset with complete coverage of all possible prompting strategies. A user study could be valuable for further characterizing the range of prompts which could be employed in practice to produce AI-generated peer reviews. However, we note that recruiting participants for such a study presents ethical challenges because our main focus is specifically on detecting unethical applications of LLMs in the peer review process. Therefore, the target population for participants in this study would be individuals who are willing to (1) admit that they engage in unethical behavior that violates the review policies of most major conferences, and (2) are willing to share the prompts that they use to engage in such behavior with researchers who are actively working to detect it. Addressing these challenges in characterizing how humans are utilizing LLMs to write peer reviews remains a valuable direction of study for future work.
Appendix H Artifact Use Consistent With Intended Use
Appendix I Reproducibility Statement
The full dataset is publicly released. All baselines are implemented using public codebases (IMGTB), and our Anchor method and prompting setup are described in detail in the main text and appendix.
Appendix J Use of AI Tool